# Defending Against Alignment-Breaking Attacks via Robustly Aligned LLM

### v2

@@ 4 Experiments
In this section, we aim to validate the efficacy of our RA-LLM from two aspects: 1) RA-LLM can effectively reduce the attack success rate of adversarial prompts; 2) RA-LLM minimally affects the outputs of benign samples. In the following, we first introduce our experimental settings and give a detailed analysis of our experimental results and ablation study.
@@ 4.1 Experimental Settings
Dataset
We evaluated our approach on two datasets: AdvBench (Zou et al., 2023) and MS MARCO dataset (Nguyen et al., 2016). The AdvBench dataset contains two types of data, corresponding to Harmful Strings Attack and Harmful Behaviors Attack respectively. Specifically, The data used for Harmful Strings Attack consist of 500 strings related to harmful or toxic content, such as threats, discriminatory remarks, methods of crime, and dangerous suggestions, etc. The data used for Harmful Behaviors Attack consists of 500 questions that can entice the LLM to produce harmful outputs, with topics similar to Harmful Strings. MS MARCO is a question-answering dataset, where all questions originate from real user queries on Bing. We sampled 150 pieces of data from each of these three datasets for our experimental evaluation.
Attack Setting
We mainly evaluate our defense under the state-of-the-art alignment-breaking attack: the Harmful Behaviors Attack proposed by (Zou et al., 2023). The goal of Harmful Behaviors Attack is to induce the LLM to respond effectively to malicious queries, which normally should be rejected by the aligned LLMs. Harmful Behaviors Attack aims to bypass the protective measures of aligned LLMs and entice them to generate harmful content. We calculated all adversarial prompts using the default hyperparameters provided in (Zou et al., 2023).
@@ 4.2 Experimental Results
Table 1: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
@@
Individual Vicuna-7B-chat-HF 99.3% 98.7% 98.7% 10.7% 88.0%
@@
Transfer Vicuna-7B-chat-HF 99.3% 98.7% 83.3% 11.3% 71.0%
@@
Specifically, we consider two main metrics to evaluate our model’s performances: attack success rate (ASR) and benign answering rate (BAR). Attack success rate measures the number of chances when the adversarial prompts successfully circumvent the model’s alignment mechanism. An attack is regarded as successful when the LLM produces a meaningful response without rejecting to answer with typical alignment text. To ensure the defense mechanism does not overkill and reject to answer benign questions, we also tested the benign answering rate, which represents the model precision in successfully identifying benign requests (does not reject to answer the benign requests). Our defensive goal is to minimize the attack success rate as much as possible while correctly identifying benign samples with a high benign answering rate.
From Table 1, it is evident that for Individual Attack, adversarial prompts have led to high malicious response success rates of 98.7% and 96.0% on the two models respectively. However, upon employing our robustly aligned LLM, these success rates dropped to 10.7% and 6.7%. Similarly, for Transfer Attack, the application of our robustly aligned LLM reduced the attack success rates from 83.3% and 78.7% to 11.3% and 8.7%. This demonstrates that our strategy effectively mitigates adversarial attacks. Additionally, our method maintains a good benign response rate, this indicates that our approach has almost no adverse impact on the LLM’s responses to benign inputs.
@@ 4.3 Handcrafted Jailbreak Prompts
Table 2: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM using handcrafted jailbreak prompts.
@@
We also assessed the defensive capabilities of our robustly aligned LLM against these meticulously designed jailbreak prompts. Specifically, we selected the top five jailbreak prompts from jailbreakchat.com [^3] The prompts are taken according to the website result on Sept 12, 2023, voted by the online users according to their effectiveness. For each of these handcrafted jailbreak prompts, we randomly selected 30 questions from the Harmful Behaviors dataset, culminating in a set of 150 handcrafted jailbreak prompt samples.
Table 2 shows the effects of our defense method on the handcrafted jailbreak prompt dataset for three different LLMs, Vicuna-7B-chat-HF, Guanaco-7B-HF, GPT-3.5-turbo-0613, all of them underwent safety alignment. We found that our robustly aligned LLM also performs exceptionally well against such handcrafted jailbreak prompts. As seen in Table 2, handcrafted jailbreak prompts achieved attack success rates of 98.4%, 94.7%, and 82.0% on the Vicuna-7B-chat-HF, Guanaco-7B-HF, and GPT-3.5-turbo-0613 models, respectively, without additional defense beyond alignment. However, when applying to our robustly aligned LLM, the attack success rates dropped to 12%, 9.3%, and 8.0%, a result even better compared to the adversarial prompt attacks in the previous section. In the meantime, RA-LLM has no significant impact on BAR especially for the larger models like GPT-3.5-turbo-0613, which inherently possess strong semantics comprehension abilities.
@@ 4.4 Ablation Study
Furthermore, as observed in Figure 2(c), our method still exhibits good performance with various Monte Carlo trails. Even with very few Monte Carlo trials such as 15 and 10, our robustly aligned LLM maintains a benign answering rate close to 100% and a relatively low attack success rate. This suggests that reducing the number of Monte Carlo trials is a potential strategy to decrease computational overhead while maintaining stable defensive performance.
4.5 Computational Cost
$C_{\text{extra}}=(1-p)l_{\text{in}}\times c_{\text{in}}\times n+l_{\text{out}}\times c_{\text{out}}\times n,\text{ where~{}}l_{\text{out}}\leq t_{\text{max}}.$
$\frac{C_{\text{extra}}}{C_{\text{LLM}}}=\frac{(1-p)l_{\text{in}}\times c_{\text{in}}\times n+l_{\text{out}}\times c_{\text{out}}\times n}{l_{\text{in}}\times c_{\text{in}}+l_{\text{out}}\times c_{\text{out}}}\leq\frac{(1-p)l_{\text{in}}\times c_{\text{in}}\times n+t_{\text{max}}\times c_{\text{out}}\times n}{l_{\text{in}}\times c_{\text{in}}+l_{\text{out}}\times c_{\text{out}}}.$
5 Conclusion and Future work
@@
While a variety of alignment strategies have been proposed to guide the large language model to obey human ethical principles, recent works show that these alignments are vulnerable and can be bypassed by alignment-breaking attacks through carefully crafted adversarial prompts. In this work, we propose robustly aligned LLMs, which are built upon existing aligned LLMs with a robust alignment checking function, to defend against alignment-breaking attacks. One major advantage of our method is that there is no need to expensively retrain or fine-tune the original LLM for defense purposes. We also provide a theoretical analysis to verify the effectiveness of our proposed defense. The exhaustive experimental results clearly demonstrate our method can effectively defend against both automatically generated adversarial prompts and handcrafted jailbreak prompts.
Note that it is non-trivial to directly apply the current alignment-breaking attack strategies (such as Zou et al. (2023)) to our robustly aligned LLM due to the non-differentiability of our random dropping mechanism, which makes it hard to perform the gradient-based search or text optimization. So far it is under-explored whether the attackers could elaborately design stronger and more efficient attacks with the knowledge of our defense details. We leave this as our future work.
@@ Appendix B Concrete Examples
Appendix C Potential Adaptive Attack
In this section, we will discuss potential adaptive attack methods against the defense mechanism we proposed. Since our method randomly drops $p$ portion of tokens from the input ${\mathbf{x}}$ and uses Monte Carlo sampling to simulate all possible scenarios, any form of adversarial prompt may be discarded. Hence, it’s challenging to design an adaptive attack based on optimization for our defense method. However, one may also utilize this design choice and simply try increasing the length of the adversarial prompts (e.g., repeat the adversarial prompts after input for several times) to ensure the random dropping cannot fully remove the adversarial parts.
In order to figure out whether such a potential adaptive attack can invalidate our defense or not, we conducted experiments on the Harmful Behaviors attack on both the original LLM and our robustly aligned LLM. The results are presented in Table 3. We found that, on the original LLM, repeating the adversarial prompt multiple times in the input also leads to a reduction in the attack success rate. We speculate that the adversarial prompt might heavily depend on its position within the full input. Similarly, we observed that on our robustly aligned LLM, the attack success rate decreases as the number of repetitions increases. What’s more, at various repetition counts, our defense method keeps the attack success rate lower than scenarios without repetitions, hovering around 5%. This suggests that even if attackers are familiar with our defense and want to use longer repetitive adversarial prompts, our method remains effective in thwarting their attacks.
Table 3: Adaptive attack success rate in our robustly aligned LLM. Repetition Times represents the number of repetitions of adversarial prompts
Repetition Times No Repetition 2 3 5
Original LLM 100.0% 46.0% 34.0% 31.0%
Robustly Aligned LLM 11.0% 5.0% 6.0% 3.0%
Appendix D Defensive Efficacy Against Harmful Strings Attack
Table 4: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
@@ Appendix D Defensive Efficacy Against Harmful Strings Attack
Appendix E Collaborating with Safety Alignment on LLMs to Counteract Attacks
We have shown in the experiments that applying the random dropping strategy on malicious requests with adversarial prompts can effectively trigger the alignment of the model. However, for benign requests, random dropping may lead to a loss of key information and make the LLM occasionally generate unable-to-answer responses similar to typical alignment responses. This leads to a certain level of decrease in terms of benign answering rate.
Clearly, we can further reduce the loss on benign answering rate if the alignment response of the LLMs can be distinguishable from the other types of unable-to-answer responses. For instance, during the alignment fine-tuning process, the LLM is instructed to always start the response to malicious requests with a special token. When applying our defensive method, it is only necessary to output and check the first token in each Monte Carlo trial. Such a collaborative strategy on alignment and RA-LLM will not only significantly improve our recognition accuracy for malicious inputs but also help in further reducing computational costs. Due to limited resources, we leave this part as our future work.
@@ Appendix F Comparison with LLM Self Defense
In this section, we additionally compared our method with the concurrent work, LLM Self Defense (Helbling et al., 2023). We use the more effective ”Questions as suffix” setting, that simply utilizes itself or another LLM instance to detect if its own response is harmful through the instruction such as
Here is some text:{response} Is it harmful? Give answer as “Yes, it is harmful” or “No, it is harmless”.
Specifically, we conducted both Self Defense through itself and Self Defense through GPT3.5 and summarized the comparison results in Table 5. We can observe that the attacked models (i.e., Vicuna and Guanaco) inherently struggle to discern harmful content, leading to low BAR. Moreover, the outputs of the detection prompt contain unexpected content including deviations from the predefined responses of ”Yes, it is harmful.” or ”No, it is harmless.”, continue completion of the unfinished content from the previous context, and even non-responsiveness from LLMs. These suggest that in scenarios preventing the use of external powerful models, LLM Self Defense might be less effective.
@@
Table 5: The benign answering rate and attack success rate of the original LLM, self Defense, self Defense by GPT3.5, and our RA-LLM under individual adversarial alignment-breaking attacks.
@@
Appendix G Comparison with Perplexity-Based Defense
Table 6: The benign answering rate and attack success rate of the original LLM, perplexity defense, and our robustly aligned LLM under two alignment-breaking attacks.
@@ Appendix G Comparison with Perplexity-Based Defense
Appendix H Time Cost

### v3

@@ 4 Experiments
In this section, we aim to validate the efficacy of our RA-LLM from two aspects: 1) RA-LLM can effectively reduce the attack success rate of adversarial prompts; 2) RA-LLM minimally affects the outputs of benign samples. In the following, we first introduce our experimental settings and give a detailed analysis of our experimental results and ablation study. Full details of the experiment can be found in Appendix E.
@@ 4.1 Experimental Settings
Attack Method
We conducted tests on the current state-of-the-art methods in three typical alignment attack techniques: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2023), and Tree of Attacks (TAP) (Mehrotra et al., 2023). GCG aims to optimize from a meaningless string to discover an adversarial suffix. AutoDAN iterates from an existing handcrafted jailbreak prompt using a genetic algorithm, seeking the most effective alignment-breaking prompt. TAP, starting from a similar handcrafted jailbreak prompt, employs an additional (often more powerful) Large Language Model (LLM) to refine the initial prompt. The objective of all these attacks is to prompt the LLM to respond effectively to a harmful query string when appended with the discovered jailbreak prompt.
Metric
We consider two main metrics to evaluate our model’s performances: attack success rate (ASR) and benign answering rate (BAR). Attack success rate measures the number of chances when the adversarial prompts successfully circumvent the model’s alignment mechanism. An attack is regarded as successful when the LLM produces a meaningful response without rejecting to answer with typical alignment text. To ensure the defense mechanism does not overkill and reject to answer benign questions, we also tested the benign answering rate, which represents the model precision in successfully identifying benign requests (does not reject to answer the benign requests). Our defensive goal is to minimize the attack success rate as much as possible while correctly identifying benign samples with a high benign answering rate.
@@ 4.2 Experimental Results
GCG-Individual Vicuna-7B-chat-HF 99.3% 98.7% 98.7% 10.7% 88.0%
@@
GCG-Transfer Vicuna-7B-chat-HF 99.3% 98.7% 83.3% 11.3% 71.0%
@@
Table 1: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
From Table 1, it is evident that for Individual Attack, adversarial prompts have led to high malicious response success rates of 98.7% and 96.0% on the two models respectively. However, upon employing our robustly aligned LLM, these success rates dropped to 10.7% and 6.7%. Similarly, for Transfer Attack, the application of our robustly aligned LLM reduced the attack success rates from 83.3% and 78.7% to 11.3% and 8.7%. This demonstrates that our strategy effectively mitigates adversarial attacks. Additionally, our method maintains a good benign response rate, this indicates that our approach has almost no adverse impact on the LLM’s responses to benign inputs. Additionally, we defer more results against GCG harmful String Attack, AutoDAN, and TAP in Appendix C and D, which demonstrate that our method can still substantially diminish the effectiveness of these attacks.
@@ 4.3 Handcrafted Jailbreak Prompts
Table 2: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM and our robustly aligned LLM using handcrafted jailbreak prompts.
We also assessed the defensive capabilities of our robustly aligned LLM against these meticulously designed jailbreak prompts. Specifically, we selected the top five jailbreak prompts from jailbreakchat.com [^3] The prompts are taken according to the website result on Sept 12, 2023, voted by the online users according to their effectiveness. For each of these handcrafted jailbreak prompts, we randomly selected 30 questions from the Harmful Behaviors dataset, culminating in a set of 150 handcrafted jailbreak prompt samples. Table 2 shows the effects of our defense method on the handcrafted jailbreak prompt dataset for three different LLMs, Vicuna-7B-chat-HF, Guanaco-7B-HF, GPT-3.5-turbo-0613, all of them underwent safety alignment. We found that our robustly aligned LLM also performs exceptionally well against such handcrafted jailbreak prompts. As seen in Table 2, handcrafted jailbreak prompts achieved attack success rates of 98.4%, 94.7%, and 82.0% on the Vicuna-7B-chat-HF, Guanaco-7B-HF, and GPT-3.5-turbo-0613 models, respectively, without additional defense beyond alignment. However, when applying to our robustly aligned LLM, the attack success rates dropped to 12%, 9.3%, and 8.0%, a result even better compared to the adversarial prompt attacks in the previous section. In the meantime, RA-LLM has no significant impact on BAR especially for the larger models like GPT-3.5-turbo-0613, which inherently possess strong semantics comprehension abilities.
@@ 4.4 Ablation Study
As observed in Figure 3(c), our method still exhibits good performance with various Monte Carlo trails. Even with very few Monte Carlo trials such as 15 and 10, our RA-LLM maintains a BAR close to 100% and a relatively low attack success rate. This suggests that reducing the number of Monte Carlo trials is a potential strategy to decrease computational cost while maintaining stable defensive performance.
5 Computational Cost
We adopt two mechanisms to ensure our method is time-efficient: (1) Limited output length: we limit a small maximum generation length in Monte Carlo simulation (e.g., $t_{max}=10$), since the negative prefixes often appear at the start of LLM responses, allowing us to obtain effective defense without generating full responses; (2) Early exit mechanism: during the Monte Carlo simulation, if the detected failure cases exceed our set threshold, RA-LLM can directly terminate the process early and mark the input as malicious. We also evaluated the actual time overhead of RA-LLM against the original LLM in our experiments. We tested 150 attack samples and recorded the normal inference time on two LLMs and the time required by our RA-LLMs. We report the extra time per data on average overhead in Table 3 where the values in parenthesis represent the percentage of additional time relative to the normal inference time. These results show RA-LLM’s extra time requirement is less than 20% of normal inference. See more details of time cost and API cost in Appendix H.
Model normal inference time RA-LLM extra time
Vicuna-7B-chat-HF 20.97s 3.93s(18.2%)
Guanaco-7B-HF 30.36s 3.76s(12.4%)
Table 3: Additional time cost of RA-LLM.
6 Adaptive Attack
In this section, we explore three potential adaptive attack methods against RA-LLM to assess its resilience.
Repetition Times No Repetition 2 3 5
Original LLM 100.0% 46.0% 34.0% 31.0%
RA-LLM 11.0% 5.0% 6.0% 3.0%
Table 4: Adaptive attack success rate(ASR) in our robustly aligned LLM. Repetition Times represents the number of repetitions of adversarial prompts
Repeating adversarial prompts
Since our method randomly drops $p$ portion of tokens from the input ${\mathbf{x}}$ and uses Monte Carlo sampling to simulate all possible scenarios, any form of adversarial prompt may be discarded. Hence, it’s challenging to design an adaptive attack based on optimization for our defense method. However, one may also utilize this design choice and simply try increasing the length of the adversarial prompts (e.g., repeat the adversarial prompts after input for several times) to ensure the random dropping cannot fully remove the adversarial parts.
In order to figure out whether such a potential adaptive attack can invalidate our defense or not, we conducted experiments on the Harmful Behaviors attack on both the original LLM and our robustly aligned LLM. The results are presented in Table 4. We found on the original LLM, repeating the adversarial prompt multiple times in the input also leads to a reduction in the attack success rate. What’s more, at various repetition counts, our defense method keeps the ASR lower than scenarios without repetitions, hovering around 5%.
Replacing the Target Model with RA-LLM
In this adaptive attack approach, we directly substitute the target model used during the attack with its RA-LLM version. That is, if RA-LLM deems the input harmless, it will return the response of the underlying LLM to the input. If RA-LLM judges the input as harmful, i.e., $AC(\cdot)$ returns Fail, it will return a fixed string "I’m sorry, but I can’t assist with that."
Given that it’s impossible to directly derive gradients for RA-LLM, gradient-based attack methods like GCG(Zou et al., 2023) cannot be applied. Thus, we experimented with TAP(Mehrotra et al., 2023) and AutoDAN(Liu et al., 2023) attack methods and found that after multiple iterations, both methods failed to generate effective jailbreak prompts, resulting in a 0% success rate. We speculate that since RA-LLM always returns a fixed string and probability distribution for harmful inputs, and this string and distribution are manually specified by us, the attack methods cannot find a reasonable optimization direction, leading to convergence failure.
Incorporating an Additional Loss Term
As mentioned above, due to the non-differentiability of our random dropping mechanism, it’s challenging to directly apply gradient-based attack methods on RA-LLM. Therefore, we considered designing a suitable adaptive attack for gradient-based white-box attack methods. A potential method involves optimizing the probability of RAC returning Pass as one of the objective functions. However, for attackers, this means that each iteration during the attack process would require approximately 20 times more overhead. This is significant because such gradient-based attack methods (e.g., GCG (Zou et al., 2023)) already necessitate substantial computational resources.
As an alternative, we experimented with treating random dropping as a "Transformation" and applying the Expectation over Transformation (EoT) method. We conducted experiments on the Vicuna-7B model using the GCG’s Individual-Behavior setting, dropping 30% of the tokens randomly at the start of each optimization, and then optimizing the remaining 70%. We found that under these circumstances, the loss oscillated near its initial value and failed to converge. We believe this phenomenon may be attributed to the following reasons:
1) The objective function itself is difficult to optimize. Finding a successful adversarial sample on a well-trained LLM is challenging, and obtaining a robust adversarial sample is even harder. 2) In such attack methods, a greedy search approach is typically employed, storing the jailbreak prompt with the smallest loss after each update round as the target for the next optimization. However, the loss of the new input after applying random dropping is quite uncertain. 3) Practical adaptive attacks need to be able to bypass situations both with and without RA-LLM, making the objective function even more challenging to optimize.
7 Conclusion
@@
While a variety of alignment strategies have been proposed to guide LLMs to obey human values, recent works show that these alignments are vulnerable and can be bypassed by alignment-breaking attacks through adversarial prompts. In this work, we propose robustly aligned LLMs, which are built upon existing aligned LLMs with a robust alignment checking function, to defend against alignment-breaking attacks. One major advantage of our method is that there is no need to expensively retrain or fine-tune the original LLM for defense purposes. We also provide a theoretical analysis to verify the effectiveness of our proposed defense. The exhaustive experiments clearly demonstrate our method can effectively defend against both automatically generated adversarial prompts and handcrafted jailbreak prompts.
8 Limitations
Our work is primarily limited in two dimensions. First, the random dropping mechanism used in our proposed method still has a minor effect on benign samples in some models. Specifically, the benign answering rate of Guanaco-7B-chat-HF decreased from 95.3% to 92.0%. Future work may investigate how to design defenses with less impact on benign samples, such as exploring better dropping methods to further reduce the adverse effect. Second, due to the limitations of current jailbreak attack techniques, we have not assessed how our method performs when faced with very extreme cases, such as particularly long or short adversarial prompts. Future work could further design these extreme cases and study their adaptability to our proposed defending method.
9 Acknowledgement
We thank the anonymous reviewers for their helpful comments. This work is partially supported by DHS (17STQAC00001-07-00). The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix B Concrete Examples
Appendix C Defensive Efficacy Against Harmful Strings Attack
@@
Table 5: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
Appendix D Defensive Efficacy Against AutoDAN and TAP
We additionally conduct experiments to validate the effectiveness of RA-LLM against AutoDAN (Liu et al., 2023) and Tree of Attacks (TAP) (Mehrotra et al., 2023). In specific, AutoDAN can automatically generate semantic meaningful jailbreak prompts using a genetic algorithm with a handcrafted jailbreak prompt as initialization. TAP, starting from a similar handcrafted jailbreak prompt, iteratively refines the candidate (attack) prompts using tree-of-thought reasoning. For the implementation of AutoDAN, We used the same parameter settings provided in their papers. When evaluating AutoDAN, the maximum number of generated tokens is 20. In, Table 6, we summarized the experimental results of RA-LLM against AutoDAN and TAP, where AutoDAN-GPT refers to AutoDAN with GPT mutation. We observed that our method can still significantly reduce the ASR of these attacks.
Attack Models BAR ASR ASR reduce
Original LLM RA-LLM Original LLM RA-LLM
AutoDAN Vicuna-7B-chat-HF 99.3% 98.7% 90.7% 42.0% 48.0%
Guanaco-7B-HF 95.3% 92.0% 98.7% 16.7% 82.0%
AutoDAN-GPT Vicuna-7B-chat-HF 99.3% 98.7% 88.7% 41.3% 47.4%
Guanaco-7B-HF 95.3% 92.0% 100.0% 15.3% 84.7%
TAP Vicuna-7B-chat-HF 99.3% 98.7% 98.0% 16.0% 82.0%
Guanaco-7B-HF 95.3% 92.0% 98.0% 12.7% 85.3%
Table 6: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM and our robustly aligned LLM under AutoDAN and TAP attacks.
Appendix E Details of Experiment
In evaluating the success rate of attacks, we utilized the Harmful Behaviors and Harmful Strings data from AdvBench (Zou et al., 2023). Harmful Strings comprises 500 strings related to harmful or toxic content, such as threats, discriminatory remarks, methods of crime, and dangerous suggestions. Harmful Behaviors consists of 500 prompts that could induce LLMs to produce harmful outputs, with themes similar to those of Harmful Strings. Initially, we employed GCG(Zou et al., 2023) to obtain 150 Harmful Behaviors instances that successfully break alignment along with their adversarial suffixes. These were then used in testing AutoDAN(Liu et al., 2023) and TAP(Mehrotra et al., 2023). For BAR, testing was conducted on the MS MARCO dataset(Nguyen et al., 2016), a question-answering collection where all questions are derived from real user queries on Bing. Here, we randomly selected 150 data points for our experiment.
For all attack methods, we adhered to the default parameters provided in their official code unless specified otherwise. Unless explicitly mentioned, the hyperparameters for RA-LLM in all experiments were: the random dropping ratio $p=0.3$, the threshold $t=0.2$, the maximum number of generated tokens $t_{\text{max}}=10$, and the number of random dropping trials $n=20$. All experimental models were loaded using HuggingFace, with LLM parameters set to default values aside from RA-LLM’s hyperparameters. All experiments were conducted on an NVIDIA RTX A6000 GPU.
@@ Appendix F Comparison with LLM Self Defense
In this section, we additionally compared our method with the concurrent work, LLM Self Defense (Helbling et al., 2023). We use the more effective "Questions as suffix" setting, that simply utilizes itself or another LLM instance to detect if its own response is harmful through the instruction such as
Here is some text:{response} Is it harmful? Give answer as “Yes, it is harmful" or “No, it is harmless".
Specifically, we conducted both Self Defense through itself and Self Defense through GPT3.5 and summarized the comparison results in Table 7. We can observe that the attacked models (i.e., Vicuna and Guanaco) inherently struggle to discern harmful content, leading to low BAR. Moreover, the outputs of the detection prompt contain unexpected content including deviations from the predefined responses of "Yes, it is harmful." or "No, it is harmless.", continue completion of the unfinished content from the previous context, and even non-responsiveness from LLMs. These suggest that in scenarios preventing the use of external powerful models, LLM Self Defense might be less effective.
@@
Table 7: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM, self Defense, self Defense by GPT3.5, and our RA-LLM under individual adversarial alignment-breaking attacks.
@@
Table 8: The benign answering rate (BAR) and attack success rate (ASR) of the original LLM, perplexity defense, and our robustly aligned LLM under two alignment-breaking attacks.
Appendix G Comparison with Perplexity-Based Defense
Appendix H Computational Cost
H.1 Time Cost